40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 16 21:13:07 2012, Suresh, Update, Electronics, REFL165 repair: PD replaced, DC response checked with a torch light PC_30641_old.jpgPD_30641_new.jpgIMG_0548.JPG
    Reply  Tue Feb 28 16:31:08 2012, Suresh, Update, Electronics, REFL165 repair: Characterization REFL165_Characterization.pdfREFL165_response_shift.pdf
       Reply  Wed Feb 29 00:22:35 2012, Suresh, Update, Electronics, REFL165 repair: Installed on the AS table 
       Reply  Wed Feb 29 01:14:40 2012, Suresh, Update, Electronics, REFL165 repair: Characterization 
Message ID: 6333     Entry time: Tue Feb 28 16:31:08 2012     In reply to: 6290     Reply to this: 6337   6339
Author: Suresh 
Type: Update 
Category: Electronics 
Subject: REFL165 repair: Characterization 

The transfer function and current noise were measured.  The location of the peak shifts with the amount of incident light power (RF or DC).  The TF was measured at an incident 1064nm light power of 0.4 mW which produced a DC output voltage of 14 mV => DC photocurrent of 0.28 mA. 

Many of the effects that Koji noted in the previous characterization are still present.

In addition I observed a shift of the peak towards lower frequencies as the RF power supplied to the AM Laser (Jenne Laser) is increased.  This could create a dependance of the demodulation phase on incident RF power.

The plots are attached below.

Attachment 1: REFL165_Characterization.pdf  272 kB  Uploaded Wed Feb 29 01:15:55 2012  | Hide | Hide all
REFL165_Characterization.pdf REFL165_Characterization.pdf REFL165_Characterization.pdf REFL165_Characterization.pdf
Attachment 2: REFL165_response_shift.pdf  34 kB  Uploaded Wed Feb 29 01:22:46 2012  | Hide | Hide all
ELOG V3.1.3-